Airline Accessories has the following current assets: cash, $95 million; receivables, $87 million; inventory, $175 million; and other current assets, $11 million. Airline Accessories has the following liabilities: accounts payable, $84 million; current portion of long-term debt, $28 million; and long-term debt, $16 million. Based on these amounts, calculate the current ratio and the acid-test ratio for Airline Accessories. (Enter your answers in millions, not in dollars. For example, $5,500,000 should be entered as 5.5.)

Current Ratio Acid-Test Ratio
Answer #1

Current Ratio = current assets/current liabilities

= (95 + 87 + 175 + 11)/(84 + 28)

= 368/112

= 3.285 = 3.29 : 1

AcidTest Ratio

= (current assets - inventory)/current liabilities

= (368 - 175)/112

= 193/112

= 1.723 = 1.72 : 1